eBook Corporate Tax Liability The Wendt Corporation reported $35 million of taxable income. Its federal tax rate was 21% (ignore any possible state corporate taxes). What is the company's federal income tax bill for the year? Enter your answer in dollars. For example, an answer of $1.2 million should be entered as 1,200,000. Round your answer to the nearest dollar. $ 7350000 Assume the firm receives an additional $1 million of interest income from some bonds it owns. What is the additional tax on this interest income? Enter your answer in dollars. For example, an answer of $1.2 million should be entered as 1,200,000. Round your answer to the nearest dollar. $ 210000 Now assume that Wendt does not receive the interest income but does receive an additional $1 million as dividends on some stock it owns. Recall that 50% of dividends received are tax exempt. What is the additional tax on this dividend income? Enter your answer in dollars. For example, an answer of $1.2 million should be entered as 1,200,000. Round your answer to the nearest dollar. $
Answer #1

Wendt Corporation:

Taxable income $35,000,000.

Federal tax rate = 21%

Tax Liability = $35,000,000 * 21% = $7,350,000

Additional Interest Income of $1,000,000

Federal tax rate = 21%

Additional Tax Liability on Interest Income = $1,000,000 * 21% = $210,000

If Additional Dividend Income of $1,000,000

Federal tax rate = 21%

Additional Tax Liability on Dividend Income = $1,000,000 * 21% = $210,000

If 50% of Dividend Income is exempt of Total Dividend Income of $1,000,000.

Federal tax rate = 21%

Additional Tax Liability on Dividend Income = $1,000,000 *50%* 21% = $105,000